Project

General

Profile

Meeting #15703

Meeting Notes (running)

Added by Jonathan Edelen over 3 years ago. Updated over 2 years ago.

Status:
New
Priority:
Normal
Assignee:
-
Start date:
03/01/2017
Due date:
% Done:

0%

Estimated time:
Duration:

Description

This will have the weekly meeting notes

History

#1 Updated by Jonathan Edelen over 3 years ago

Meeting Notes 1 March 2017

  • An issue with the main injector began about a week ago
    • beam loss at high field under high average current conditions
    • When the reached the highest intensity the beam was getting pushed into the outside wall
  • Digital radial position system then diverged from the analog radial position system as well as the Q530 BPM
  • Placed two high power 3db pads at the end of the BPM signal cables which seems to have fixed the issue
  • Possibly a thermal effect due to higher power operations

#2 Updated by Jonathan Edelen over 3 years ago

Meeting notes 15 March 2017

  • No new complaints on the changes made two weeks ago

#3 Updated by Jonathan Edelen over 3 years ago

Meeting notes 22 March 2017

  • No new information, this needs study

#4 Updated by Jonathan Edelen over 3 years ago

Meeting notes 29 March 2017

  • Nothing new, this needs study
  • Philip will follow up on a down day

#5 Updated by Jonathan Edelen over 3 years ago

Meeting Notes 5 April 2017

  • Need to follow up on a down day.
  • Issue with the recycler came back when the new boards were installed, but is now resolved. There needs to be some new code installed.

#6 Updated by Jonathan Edelen over 3 years ago

Meeting notes 12 April 2017

  • Still need to follow up on a down day...

#7 Updated by Jonathan Edelen over 3 years ago

Notes 19 April 2017

  • New problem has been reported
    • Stray F-3 events being broadcast that are sourced from our recycler crate
    • This has become more frequent
    • Seems to be happening during end cycle
    • Philip will investigate

#8 Updated by Jonathan Edelen over 3 years ago

  • Tracker changed from Bug to Meeting

Meeting notes 26 April 2017

  • Working on some changes to the MRF portion as requested by John
    • This is to distribute the load when compensating for a failed cavity
  • Stray F-3 event
    • We have confirmed that it is being generated in our system
    • Continuing to investigate

#9 Updated by Jonathan Edelen over 3 years ago

Meeting Notes 3 May 2017

  • Narrowed the problem down to 4 lines of assembly language
  • Still not resolved

#10 Updated by Jonathan Edelen over 3 years ago

Meeting notes 10 May 2017

  • Attempted diagnostics masked the problem
  • Tests in different configurations (some worked for short periods some didn't)
  • We think the problem is narrowed down
    • Writes from the DSP to the FPGA, there appears to be a glitch
    • Appears to be a timing issue (writing without handshakes)
    • Philip is working on a solution

#11 Updated by Jonathan Edelen over 3 years ago

Meeting Notes 17 May 2017

  • The issue was fixed last week, tested in the lab, and put in the operational machine during the shutdown yesterday
  • We should continue to monitor (controls keeps track of the excess events)
  • There is a detailed report on redmine
  • There was a call regarding the UCD not decoding the event
    • Possibly do some parasitic testing while MI is shutdown before the recycler is shutdown
  • We will do an upgrade of the hardware

#12 Updated by Jonathan Edelen over 3 years ago

Meeting notes 24 May 2017

  • In a holding pattern for the time being due to hardware issues

#13 Updated by Jonathan Edelen over 3 years ago

Meeting notes 7 June 2017

  • Testing new firmware in the lab for the UCD modules
    • Testing of this occurred in the recycler last year (successful)
  • New slot-0 controller means new libraries
  • Would like to test before the shutdown
  • Wanted to shrink the timing by 15 milliseconds for the transfer
  • More mysterious behavior

#14 Updated by Jonathan Edelen over 3 years ago

Meeting notes 12 July 2017

  • MI and Recycler are now shut down until October
  • The events and timelines are available so we should be able to test our systems
  • There was an issue with locking to the marker, which has been resolved (error in the firmware that was introduced when changing SOM modules)
  • Need to be a little better about fully testing changes that are going into operational systems
  • Look at the comb filter system

#15 Updated by Jonathan Edelen about 3 years ago

Meeting notes 2 Aug 2017

  • Note we are currently in a shutdown
  • There is some hardware being installed by the high level group
    • Use three recycler cavities
    • One could be either an A or B cavity, modules for switching the fanback signals
    • It is transparent to our system at the moment but that may change
  • Some changes in MI and MRF that were requested a little while ago
  • Currently being tested then try out at MI-60 in the future

#16 Updated by Jonathan Edelen about 3 years ago

Meeting notes 16 Aug 2017

  • Nothing new to report

#17 Updated by Jonathan Edelen about 3 years ago

Meeting notes 30 Aug 2017

  • Made some changes to the code that will implement single group operation. Currently in lab testing
    • Exploring options for switching back and forth
    • Code will be commented and documented
  • Need to talk with Duane about the libraries

#18 Updated by Jonathan Edelen about 3 years ago

Meeting notes 20 September 2017

  • Shutdown continues
  • Testing of new cycles that they want to run when the machine comes back on
  • Shutdown ends the week of October 23rd
  • Comb filter discussion is ongoing

#19 Updated by Joshua Einstein about 3 years ago

Meeting Notes 27 September 2017

  • Shutdown continues
  • Testing ongoing
  • Awaiting end of shutdown for operational testing.

#20 Updated by Joshua Einstein almost 3 years ago

Meeting Notes 25 October 2017

  • Shutdown continues, but turn-on is commencing.
  • No additional news.

#21 Updated by Joshua Einstein almost 3 years ago

Meeting Notes 10 November 2017

  • Ongoing bring-up. Initial issue was a power amplifier that wasn't powered on.
  • Secondary issue was due to connecter problem behind rack
  • Week following next is most likely time for final bring-up
  • Want diode detector instrumentation for reference signal loss
  • 2.5 MHz system running off SoC MFC

#22 Updated by Joshua Einstein over 2 years ago

Meeting Notes 07 February 2018

  • Checking on vxWorks updates and future plans
  • MI cycle has been shortened to the point where the end of beam is interfering with our processing time

#23 Updated by Joshua Einstein over 2 years ago

Meeting Notes 28 February 2018

  • Installed updated VXI-UCD cards with new firmware in MI and RR systems
  • 1 week run without any errors noticed
  • Need to test with injection events in timeline
  • RR ring frequency scans having issues

#24 Updated by Joshua Einstein over 2 years ago

Meeting Notes 7 March 2018

  • Still no errors after running on updated UCD code
  • Do we need Duane to bump up version of libraries. Is this possible given other systems?
    • We believe these libraries are only used in MI system
    • Possibly backup the production version locally or in production folder

#25 Updated by Joshua Einstein over 2 years ago

Meeting Notes 25 April 2017

  • Duane made production version of libraries
  • Old production version backed up

#26 Updated by Joshua Einstein over 2 years ago

Meeting Notes 2 May 2018

  • 5500 board died last week (Friday) and looked as if it had corrupted boot parameters
  • Log book entry


Also available in: Atom PDF